# A Framework for Objective-Driven Dynamical Stochastic Fields

### v1

@@ 3.3 How the Objective Operators may be Designed
We begin from the two-entity formulation ${\Omega}={\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}}\times{\mathcal{N}}$ where the objective operator is the easiest to make sense of. Let us view one entity $x_{e}$ as the “environment” and the other entity $x_{a}$ as an “agent” that interacts with the environment. We focus on the entity $x_{a}$ where it minimizes its objective value given by an objective operator ${\bm{\Gamma}}$. Note that the objective operator ${\bm{\Gamma}}$ needs to be local, i.e., it only depends on what the entity $x_{a}$ observes.
@@
Notation. We use ${\bm{\Gamma}}^{\prime}\leftarrow{\bm{\Gamma}}$ to denote that the objective signal given by ${\bm{\Gamma}}$ appears as an objective signal given by ${\bm{\Gamma}}^{\prime}$ after a finite response time. The important property we need is that they induce the same objective value, i.e., ${\bm{\Gamma}}^{\prime}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$, as a consequence of not missing propagating signals.
It is not really satisfying if the objective value can only be passed unchanged. Indeed, there is nothing to stop the outer entity $x_{1}$ from do something about it. Let us denote this “something” as a linear operator ${{\mathbf{P}}}:{\widetilde{\mathcal{H}}}\to{\widetilde{\mathcal{H}}}$, and thus
@@
Here is a crucial observation. In order for the inner and outer entities, when viewed as whole, to minimize the original objective value given by ${\bm{\Gamma}}$, the objective propagation operators ${{\mathbf{P}}}$ must be chosen carefully. Given that the inner entity is obliged to minimize its objective value ${\bar{\gamma}}^{\prime}={\bm{\Gamma}}^{\prime}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bm{\Gamma}}{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$, it should align with the original objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ (up to a constant shift). This means that we would want ${\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ to be a fixed point of the operator ${{\mathbf{P}}}$, i.e.,
${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (3.27)
Now, let us see the world from entity $x_{1}$’s perspective, where $x_{e}$ and $x_{2}$ forms its environment. Without any additional assumptions, there is no way for $x_{1}$ to tell whether $x_{2}$ is also connected to $x_{e}$. Therefore, given the symmetry, it is reasonable to expect that entity $x_{2}$ also generates objective signals to entity $x_{1}$. As a result, entity $x_{2}$ would also choose an objective propagation operator ${{\mathbf{P}}}^{\prime}$ and generate the objective signal ${\bm{\Gamma}}^{\prime\prime}$ to provide to entity $x_{1}$, i.e.,
@@
To balance the two objective signals, let us simply introduce some weights for them, and thus making the final objective operator for entity $x_{1}$ a weighted average of ${\bm{\Gamma}}$ and ${\bm{\Gamma}}^{\prime\prime}$.
@@ 3.4 Objective Propagation
$\forall x,x^{\prime}:\quad{\omega}^{\prime}_{|{{\mathcal{U}}}_{x}}={\omega}_{|{{\mathcal{U}}}_{x}}\quad\implies\quad{{\mathbf{P}}}_{xx^{\prime}}{\hat{\bm{e}}}_{{\omega}^{\prime}{\omega}}=0$ (3.29)
In designing the objective operators for the acting entities, as discussed in the previous subsection, we expect ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$, such that the propagated objective signals stays meaningful and aligned with the “true” objective signal given by the environmental entities. Observe that this would require the propagation operators ${{\mathbf{P}}}$ to depend on the field dynamics ${{\mathbf{G}}}$, except in the trivial case ${{\mathbf{P}}}=\mathbf{I}$.
Therefore, the objective operator of each acting entity is given by objective signals propagated from other entities. The propagation is given by ${\bm{\Gamma}}_{x}\leftarrow{\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x}$ where ${\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}$. For each entity, we may picture this operation as both propagating and generating objective signals to its neighbors. The final effect of objective propagation can be effectively captured by the following equations.
@@
${\bm{\Gamma}}_{x}={\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where }\ {\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x},\forall x^{\prime}\in{\mathcal{X}}_{a},$ (3.30)
@@
Note that this results in a non-local objective operator ${\bm{\Gamma}}_{x}$ serving as an effective objective operator to characterize what an entity $x$ will eventually receive given any jump, possibly non-local to $x$, in the field. This simplifies the calculation and ensures the same ${\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ as desired. Another perspective is to consider the speed of objective propagation to be much faster than the propagation of field actions, resulting in a seemingly instantaneous non-local objective propagation. For notational ease, we use the same ${\bm{\Gamma}}_{x}$ to denote both the original local objective operator and the corresponding non-local effective objective operator. However, in the rest of the paper, we will exclusively use ${\bm{\Gamma}}_{x}$ to refer to the effective objective operator that satisfies the aforementioned propagation equation.
Intuitively, the objective propagation should preserve the same objective value, given ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ and $\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}=1$. We can show that this is indeed the case, provided the acting entities are connected strongly.
@@
  • ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$  for every acting entity $x^{\prime}$ connecting to acting entity $x$;
  • ${\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bar{\gamma}}$  for every environmental entity $x^{\prime}$ connecting to acting entity $x$;
@@
$\forall x\in{\mathcal{X}}_{a}:{\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\bar{\gamma}}.$ (3.31)
However, there would be a crucial problem if we attempt to calculate the gradient of ${\bar{\gamma}}$ w.r.t. the local generators. For example, let us impose an infinitesimal variation $\delta{{\mathbf{G}}}(x)$ to the local generator at $x$. This would induce a variation $\delta{\hat{\bar{\bm{\varphi}}}}$ in the system’s stationary state ${\hat{\bar{\bm{\varphi}}}}$. Moreover, since each objective propagator ${{\mathbf{P}}}_{x}$ depends on the system dynamics, it would also incur a variation $\delta{{\mathbf{P}}}_{x}$ for all entities. Since each objective operator ${\bm{\Gamma}}_{x}$ depends on the objective propagators, there would also be a variation $\delta{\bm{\Gamma}}_{x}$ in the effective objective operator. Therefore, this results in an incurred variation in the objective value that entity $x$ receives.
$\delta{\bar{\gamma}}(x)=\delta({\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})+\delta({\bm{\Gamma}}_{x}){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}.$ (3.32)
There are two terms arising from $\delta{\bar{\gamma}}$ as shown above. The first term, ${\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}})$, can be seen as the case where the objective operators are fixed. Thus, it reduces to the scenario previously studied, and the corresponding result is essentially given by Theorem 3.20. However, the second term involves $\delta({\bm{\Gamma}}_{x})$, which requires solving the following global equation derived from the objective propagation equation (Definition 3.24):
$\delta{\bm{\Gamma}}_{x}={\Lambda}^{x^{\prime}}_{x}\delta{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where}\ \ \delta{\bm{\Gamma}}_{x^{\prime}x}=\delta{\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}+{\bm{\Gamma}}_{x^{\prime}}\delta{{\mathbf{P}}}_{x^{\prime}x}.$ (3.33)
@@ 3.5 An Interesting Choice of the Propagation Operator ${{\mathbf{P}}}[{{\mathbf{Q}}}]$
The motivation for this propagator is simple: instead of just passing the objective signal untouched, let us also convey some sorts of gradient information alongside it. It turns out, interestingly, it leads to local gradient computation.
@@
${\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\hat{\bar{\bm{\varphi}}}}.$ (3.34)
We can observe that the operator $(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}})$ appears just like how an objective propagator would appear.
@@
It turns out that ${{\mathbf{P}}}$ is a well-defined objective propagator with the following nice properties.
@@
    ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta{{\mathbf{P}}}$ satisfies $\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
However, there is a problem: the above propagator ${{\mathbf{P}}}$ is not local, meaning that it would react to any dynamics within the system. To ensure it only reacts to local dynamics, we may insert a local operator ${{\mathbf{Q}}}:{\mathcal{H}}\to{\mathcal{H}}$ to filter out distant dynamics. As we will see, this operator needs to be closed on ${\Delta\mathcal{H}}$, i.e., ${{\mathbf{Q}}}{\Delta\mathcal{H}}\subseteq{\Delta\mathcal{H}}$. In fact, such operators are easy to find. For example, any linear operator that preserves the normalization will be closed on ${\Delta\mathcal{H}}$. As shown below, this little modification preserves essentially all the desirable properties of ${{\mathbf{P}}}$.
@@
For example, a natural choice of the local operator ${{\mathbf{Q}}}_{x}$ for $x\in{\mathcal{X}}$ can be simply as taking the expectation over the stationary distribution conditioned on the local information. Formally,
${{\mathbf{Q}}}_{x}{\hat{\bm{e}}}_{{\omega}}=\left({\hat{\bm{e}}}_{{\omega}^{\prime}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}}_{x,{\omega}}{\hat{\bar{\bm{\varphi}}}}\right)/Z,$ (3.39)
where ${{\check{\bm{\eta}}}}^{{\omega}^{\prime}}_{x,{\omega}}:=\delta_{{\omega}_{|{{\mathcal{U}}}_{x}}}^{{\omega}^{\prime}_{|{{\mathcal{U}}}_{x}}}\cdot{{\check{\bm{\eta}}}}^{{\omega}^{\prime}}$ filters out the configuration disagreeing with ${\omega}$ on the local neighborhood ${{\mathcal{U}}}_{x}$, and $Z:=\sum_{{\omega}^{\prime}}{{\check{\bm{\eta}}}}^{{\omega}^{\prime}}_{x,{\omega}}{\hat{\bar{\bm{\varphi}}}}$ is a normalization factor to keep the outcome representing a probability distribution. Since ${{\mathbf{Q}}}_{x}$ preserves the normalization, it is closed on ${\Delta\mathcal{H}}$ (by Fact 3.11).
Nonetheless, any operator ${{\mathbf{Q}}}$ that is closed on ${\Delta\mathcal{H}}$ would ensure that ${{\mathbf{P}}}[{{\mathbf{Q}}}]$ has same desirable properties as ${{\mathbf{P}}}$, which we will derive following the next definition.
Definition 3.30 (Bilinear Map (,)($\cdot$,$\cdot$)( ⋅ , ⋅ )).
Given two operators ${{\mathbf{Q}}},{{\mathbf{Q}}}^{\prime}\colon{\Delta\mathcal{H}}\to{\Delta\mathcal{H}}$, we define a bilinear map (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) as the following.
@@
    ${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}={\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}])$ satisfies $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$.
Having enough tools at hand, we can show that with this choice of objective propagators, we have $\delta({\bm{\Gamma}}_{x}){\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}=0$, and thus the gradient of ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\hat{\bar{\bm{\varphi}}}}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ can be computed locally.
@@
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}{\hat{\bar{\bm{\varphi}}}}.$ (3.41)
Theorem 3.32 and Proposition 3.31 together reveal several interesting properties of the propagator ${{\mathbf{P}}}[{{\mathbf{Q}}}]$. In particular, this propagator has a similar form of the operator ${{\mathbf{P}}}$ that appears in the gradient formula, suggesting that it may encode certain gradient-related information. This connection could be potentially beneficial for estimating gradients of local entities. However, its precise implications remain unclear and are left for future work.

### v2

@@ 3.3 How the Objective Operators may be Designed
We begin with the two-entity formulation ${\Omega}={\mathcal{A}}\times{\mathcal{B}}\times{\mathcal{M}}\times{\mathcal{N}}$ where the objective operator is the easiest to make sense of. Let us view one entity $x_{e}$ as the “environment” and the other entity $x_{a}$ as an “agent” that interacts with the environment. We focus on the entity $x_{a}$ where it minimizes its objective value given by an objective operator $\bm{\Gamma}$}. Note that the objective operator $\bm{\Gamma}$} needs to be local, i.e., it only depends on what the entity $x_{a}$ observes.
@@
Notation. We use ${\bm{\Gamma}}^{\prime}\leftarrow{\bm{\Gamma}}$ to denote that the objective signal given by $\bm{\Gamma}$} appears as an objective signal given by $\bm{\Gamma}$$^{\prime}$ after a finite response time. The important property we need is that they induce the same objective value, i.e., ${\bm{\Gamma}}^{\prime}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$, as a consequence of not missing propagating signals.
It is not really satisfying if the objective value can only be passed unchanged. Indeed, there is nothing to stop the outer entity $x_{1}$ from doing something about it. Let us denote this “something” as a linear operator $\mathbf{P}$$\widetilde{\mathcal{H}}$$\to{\widetilde{\mathcal{H}}}$, and thus
@@
Here is a crucial observation. In order for the inner and outer entities, when viewed as a whole, to minimize the original objective value given by $\bm{\Gamma}$}, the objective propagation operators $\mathbf{P}$}} must be chosen carefully. Given that the inner entity is obliged to minimize its objective value ${\bar{\gamma}}^{\prime}={\bm{\Gamma}}^{\prime}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bm{\Gamma}}{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$, it should align with the original objective value ${\bar{\gamma}}={\bm{\Gamma}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ (up to a constant shift). This means that we would want ${\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ to be a fixed point of the operator $\mathbf{P}$}}, i.e.,
${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (3.27)
Now, let us see the world from entity $x_{1}$’s perspective, where $x_{e}$ and $x_{2}$ form its environment. Without any additional assumptions, there is no way for $x_{1}$ to tell whether $x_{2}$ is also connected to $x_{e}$. Therefore, given the symmetry, it is reasonable to expect that entity $x_{2}$ also generates objective signals to entity $x_{1}$. As a result, entity $x_{2}$ would also choose an objective propagation operator $\mathbf{P}$$^{\prime}$ and generate the objective signal $\bm{\Gamma}$$^{\prime\prime}$ to provide to entity $x_{1}$, i.e.,
@@
To balance the two objective signals, let us simply introduce some weights for them, thus making the final objective operator for entity $x_{1}$ a weighted average of $\bm{\Gamma}$} and $\bm{\Gamma}$$^{\prime\prime}$.
@@ 3.4 Objective Propagation
$\forall x,x^{\prime}:\quad{\omega}^{\prime}_{|{{\mathcal{U}}}_{x}}={\omega}_{|{{\mathcal{U}}}_{x}}\quad\implies\quad{{\mathbf{P}}}_{xx^{\prime}}\lvert{\omega}^{\prime}{\omega}\rangle=0$ (3.29)
In designing the objective operators for the acting entities, as discussed in the previous subsection, we expect ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$, such that the propagated objective signals stay meaningful and aligned with the “true” objective signal given by the environmental entities. Observe that this would require the propagation operators $\mathbf{P}$}} to depend on the field dynamics $\mathbf{G}$}}, except in the trivial case ${{\mathbf{P}}}=\mathbf{I}$.
Therefore, the objective operator of each acting entity is given by objective signals propagated from other entities. The propagation is given by ${\bm{\Gamma}}_{x}\leftarrow\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x}$ where ${\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}$. For each entity, we may picture this operation as both propagating and generating objective signals to its neighbors. The final effect of objective propagation can be effectively captured by the following equations.
@@
${\bm{\Gamma}}_{x}=\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where }\ {\bm{\Gamma}}_{x^{\prime}x}={\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x},\forall x^{\prime}\in{\mathcal{X}}_{a},$ (3.30)
@@
Note that this results in a non-local objective operator ${\bm{\Gamma}}_{x}$ serving as an effective objective operator to characterize what an entity $x$ will eventually receive given any jump, possibly non-local to $x$, in the field. This simplifies the calculation and ensures the same ${\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}{\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ as desired. Another perspective is to consider the speed of objective propagation to be much faster than the propagation of field actions, resulting in a seemingly instantaneous non-local objective propagation. For notational ease, we use the same ${\bm{\Gamma}}_{x}$ to denote both the original local objective operator and the corresponding non-local effective objective operator. However, in the rest of the paper, we will exclusively use ${\bm{\Gamma}}_{x}$ to refer to the effective objective operator that satisfies the aforementioned propagation equation.
Intuitively, the objective propagation should preserve the same objective value, given ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ and $\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}=1$. We can show that this is indeed the case, provided that the acting entities are connected strongly.
@@
  • ${{\mathbf{P}}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ for every acting entity $x^{\prime}$ connecting to acting entity $x$;
  • ${\bm{\Gamma}}_{x^{\prime}x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bar{\gamma}}$ for every environmental entity $x^{\prime}$ connecting to acting entity $x$;
@@
$\forall x\in{\mathcal{X}}_{a}:{\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\bar{\gamma}}.$ (3.31)
However, there would be a crucial problem if we attempt to calculate the gradient of $\bar{\gamma}$} w.r.t. the local generators. For example, let us impose an infinitesimal variation $\delta{{\mathbf{G}}}(x)$ to the local generator at $x$. This would induce a variation $\delta{\lvert\bar{\varphi}\rangle}$ in the system’s stationary state $\lvert\bar{\varphi}\rangle$}. Moreover, since each objective propagator ${{\mathbf{P}}}_{x}$ depends on the system dynamics, it would also incur a variation $\delta{{\mathbf{P}}}_{x}$ for all entities. Since each objective operator ${\bm{\Gamma}}_{x}$ depends on the objective propagators, there would also be a variation $\delta{\bm{\Gamma}}_{x}$ in the effective objective operator. Therefore, this results in an incurred variation in the objective value that entity $x$ receives.
$\delta{\bar{\gamma}}(x)=\delta({\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})={\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})+\delta({\bm{\Gamma}}_{x}){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}.$ (3.32)
There are two terms arising from $\delta{\bar{\gamma}}$ as shown above. The first term, ${\bm{\Gamma}}_{x}\delta({\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle})$, can be seen as the case where the objective operators are fixed. Thus, it reduces to the scenario previously studied, and the corresponding result is essentially given by Theorem 3.20. However, the second term involves $\delta({\bm{\Gamma}}_{x})$, which requires solving the following global equation derived from the objective propagation equation (Definition 3.24):
$\delta{\bm{\Gamma}}_{x}=\sum_{x^{\prime}}{\Lambda}^{x^{\prime}}_{x}\delta{\bm{\Gamma}}_{x^{\prime}x},\qquad\text{where}\ \ \delta{\bm{\Gamma}}_{x^{\prime}x}=\delta{\bm{\Gamma}}_{x^{\prime}}{{\mathbf{P}}}_{x^{\prime}x}+{\bm{\Gamma}}_{x^{\prime}}\delta{{\mathbf{P}}}_{x^{\prime}x}.$ (3.33)
@@ 3.5 An Interesting Choice of the Propagation Operator ${{\mathbf{P}}}[{{\mathbf{Q}}}]$
The motivation for this propagator is simple: instead of just passing the objective signal untouched, let us also convey some sort of gradient information alongside it. Interestingly, it turns out that it leads to local gradient computation.
@@
${\bm{\Gamma}}(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}}){{\widetilde{\mathbf{A}}}}(x)^{{\omega}}_{{\omega}^{\prime}}{\lvert\bar{\varphi}\rangle}.$ (3.34)
We can observe that the operator $(1+{\widetilde{\mathbf{G}}}{{\mathbf{S}}}{\mathbf{\Pi}})$ appears just as an objective propagator would.
@@
It turns out that $\mathbf{P}$}} is a well-defined objective propagator with the following desirable properties.
@@
    ${{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta{{\mathbf{P}}}$ satisfies $\delta{{\mathbf{P}}}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
However, there is a problem: the above propagator $\mathbf{P}$}} is not local, meaning that it would react to any dynamics within the system. To ensure it only reacts to local dynamics, we may insert a local operator $\mathbf{Q}$$\mathcal{H}$$\to{\mathcal{H}}$ to filter out distant dynamics. As we will see, this operator needs to be closed on $\Delta\mathcal{H}$}, i.e., ${{\mathbf{Q}}}{\Delta\mathcal{H}}\subseteq{\Delta\mathcal{H}}$. In fact, such operators are easy to find. For example, any linear operator that preserves the normalization will be closed on $\Delta\mathcal{H}$}. As shown below, this little modification preserves essentially all the desirable properties of $\mathbf{P}$}}.
@@
Any operator $\mathbf{Q}$}} that is closed on $\Delta\mathcal{H}$} would ensure that ${{\mathbf{P}}}[{{\mathbf{Q}}}]$ has the same desirable properties as $\mathbf{P}$}}, which we will derive following the next definition.
Definition 3.30 (Bilinear Map $(\cdot,\cdot)$).
Given two operators ${{\mathbf{Q}}},{{\mathbf{Q}}}^{\prime}\colon{\Delta\mathcal{H}}\to{\Delta\mathcal{H}}$, we define a bilinear map $(\cdot,\cdot)$ as the following.
@@
    ${{\mathbf{P}}}[{{\mathbf{Q}}}]{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}={\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$.
@@
    Given a variation $\delta{{\mathbf{G}}}$, the incurred $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}])$ satisfies $\delta({{\mathbf{P}}}[{{\mathbf{Q}}}]){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$.
Having enough tools at hand, we can show that with this choice of objective propagators, we have $\delta({\bm{\Gamma}}_{x}){\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}=0$, and thus the gradient of ${\bar{\gamma}}(x)={\bm{\Gamma}}_{x}{\widetilde{\mathbf{G}}}{\lvert\bar{\varphi}\rangle}$ w.r.t. local generator ${{\mathbf{G}}}(x)$ can be computed locally.
@@
$\frac{\partial{\bar{\gamma}}(x)}{\partial G(x)^{{\omega}^{\prime}}_{{\omega}}}={\bm{\Gamma}}_{x}{{\mathbf{P}}}{{\widetilde{\mathbf{A}}}}(x)_{{\omega}^{\prime}}^{{\omega}}{\lvert\bar{\varphi}\rangle}.$ (3.40)
Theorem 3.32 and Proposition 3.31 together reveal several interesting properties of the propagator ${{\mathbf{P}}}[{{\mathbf{Q}}}]$. In particular, this propagator has a similar form to the operator $\mathbf{P}$}} that appears in the gradient formula, suggesting that it may encode certain gradient-related information. This connection could potentially be beneficial for estimating the gradients of local entities. However, its precise implications remain unclear and are left for future work.